The irradiance of a beam of natural light is 400 W/m². It impinges on the first of two consecutive ideal linear polarizers whose transmission axes are 40.0° apart. How much light emerges from the two?


1
Expert's answer
2021-06-21T11:39:24-0400

"I=I_0\cos^2\alpha,"

"I=400\cdot \cos^2 40°=235~\frac{W}{m^2}."